# Decoding Neural Emotion Patterns through Large Language Model Embeddings

### v2

@@ Abstract
Linking emotional expression in natural language to brain function remains a central challenge in neuroscience and affective computing. While neuroimaging offers insight into emotion localization, it is costly and limited to controlled settings. In this study, we introduce a computational framework that maps textual emotional content to neuroanatomical brain regions, enabling scalable, imaging-free emotion–brain research. Our proposed method combines semantic embeddings, dimensionality reduction, and clustering to identify emotional states, which are mapped onto brain regions associated with affective processing. We applied this framework across three experiments: i) comparing conversational data from healthy and depressed individuals, ii) analyzing a large-scale emotion dataset, and iii) contrasting human language with outputs from a large language model. Emotional intensity was quantified using a lexical scoring system sensitive to keywords, syntax, and modifiers. This approach produced neuro-anatomically plausible mappings with high spatial specificity. Depressed subjects exhibited reduced engagement of the left insula and raphe nuclei, regions linked to interoceptive awareness and serotonergic function. The framework reliably differentiated discrete emotions and revealed that LLM-generated texts, while capturing basic affective patterns, lacked the nuanced activation seen in human language, particularly in empathy and self-referential regions such as the medial prefrontal and posterior cingulate cortices. This work establishes a scalable, cost-effective alternative to neuroimaging, advancing both clinical and computational models of emotion, while providing a neuro-inspired benchmark for assessing how closely AI-generated language mirrors human emotional expression.
@@
The integration of semantic embeddings and neuro-anatomical mapping successfully differentiated between healthy and depressed populations through distinct activation patterns.
@@
Regional assignment patterns showed strong consistency with established neuroimaging research.
In favor of reproducible research and to advance the field, all programming code/data used in this study is made publicly available.
@@ 1 Introduction
Understanding the neural correlates of emotion has traditionally relied on neuroimaging modalities such as electroencephalography (EEG) and functional magnetic resonance imaging (fMRI) [1, 2, 3]. These techniques have revealed the involvement of regions like the amygdala, insula, anterior cingulate cortex, and prefrontal cortex across different emotional states [4]. Meta-analyses of neuroimaging studies have consistently identified these key regions across diverse emotional paradigms, with the amygdala showing particular importance for threat detection and fear processing [5], while the anterior cingulate cortex and insula demonstrate critical roles in emotional awareness and interoceptive processing [6, 7]. However, traditional neuroimaging approaches face significant limitations, including high costs, restricted accessibility, controlled laboratory requirements, and limited ecological validity when studying naturalistic emotional expression [8, 9]. The increasing availability of digital text data presents unprecedented opportunities for emotion-brain mapping that could overcome these existing constraints.
Caucheteux et al. [10] demonstrated that large language model (LLM) embeddings align with human brain activity without fine-tuning, showing that pre-trained language models inherently capture aspects of neural language representations. This foundational work was extended by Toneva et al. [11], who introduced the concept of brain embeddings, highlighting the geometric alignment between the representational spaces of LLMs and brain activity during reading and listening tasks.
Further evidence of this alignment comes from Schrimpf et al. [12], who showed that artificial neural networks, especially transformer-based architectures as used in LLMs, can predict human neural responses to language with remarkable accuracy. These findings collectively support the feasibility of leveraging LLM-derived embeddings to model brain activity, suggesting that the semantic representations learned by these models may reflect fundamental aspects of how the human brain processes language and emotion.
Parallel efforts have explored mapping emotional text representations to neuro-biological substrates. Tomasino et al. [13] developed a cognitive-affective framework demonstrating how emotionally charged linguistic input recruits distinct neural systems, while Chen et al. [14] correlated sentiment analysis outputs with fMRI patterns, confirming the brain’s differentiation of emotional valence during narrative comprehension. These studies build upon earlier work demonstrating that emotional processing involves distributed neural networks, with positive emotions preferentially engaging left prefrontal regions and negative emotions showing stronger right hemisphere activation [15].
Zhou et al. [16] extended this work by associating semantic embeddings from emotional narratives with fMRI-derived brain states, particularly noting strong alignment in medial prefrontal and temporal regions. Similarly, Xiao et al. [17] applied unsupervised clustering on emotion-labeled text to uncover latent emotional dimensions and correlated them with EEG and fMRI features. These computational approaches align with neuroimaging meta-analyses showing that different emotional categories activate distinct but overlapping brain networks, with cognitive emotions recruiting prefrontal cortical areas [18, 19].
The utility of natural text has been further emphasized by Hoemann et al. [20], who highlighted the value of social media and dialogue data for studying emotion in real-world contexts. Their findings support using large-scale spontaneous language datasets to infer affective brain states, moving beyond the artificial constraints of laboratory-based emotion elicitation paradigms. This approach is particularly relevant given further research showing that naturalistic emotional expression differs significantly from laboratory-induced emotions in both linguistic patterns and associated neural activity [20].
Despite these advances, existing studies have not established a fully computational, imaging-free method that directly links textual emotional content to specific neuro-anatomical regions. Current approaches typically require either controlled laboratory settings that limit validity, or a focus on general emotional dimensions rather than specific brain region mapping [8].
This represents a significant gap in our ability to study emotional processing at scale. The motivation for developing a purely computational emotion–brain mapping framework, therefore, stems from several converging factors. First, the exponential growth of digital text data offers an unprecedented window into human emotional expression in naturalistic contexts [20]. Second, advances in semantic embedding technologies have created powerful tools for capturing nuanced relationships between language, meaning, and affective content. Third, decades of neuroimaging research have established a robust understanding of the neuro-anatomical basis of emotion processing, with meta-analytic evidence highlighting key neural circuits involved in emotional regulation and expression [18, 21].
The theoretical foundation of our proposed approach rests on the principle that emotional expression in language reflects underlying neural processes. This hypothesis is supported by evidence showing that the human brain encodes language through distributed, continuous representations [22, 23, 24]. Studies have demonstrated a direct alignment between high-dimensional language model embeddings and neural activation patterns in language-processing regions [23, 24, 25].
Recent work by Goldstein et al. [26] further established that simple linear encoding models can reliably map language model embeddings onto brain activity during natural conversation, achieving robust correlations with recorded signals. Their findings showed that deep embeddings consistently outperform traditional symbolic linguistic features (e.g., phonemes, syntactic categories) in predicting neural responses across cortical regions, suggesting that high-dimensional continuous representations more faithfully reflect the brain’s internal language mechanisms. Importantly, they also revealed individual and regional differences in neural–embedding alignment, underscoring the capacity of computational models to capture variability in neural organization that extends beyond region-specific specialization. Such evidence not only validates embedding-based approaches for brain-language mapping but also provides a methodological and theoretical foundation for extending this framework to emotion, where capturing both shared and individual patterns of neural processing is particularly critical.
Furthermore, individuals with certain mental health conditions may exhibit distinct language patterns, including variations in word choice, emotional tone, and syntactic complexity that correlate with well-documented neuro-biological abnormalities [27, 28, 8, 29, 30, 31, 32]. Extending this perspective, characteristic language patterns in clinical populations may manifest as distinct clusters within embedding space, reflecting altered cognitive and affective processing [33, 34, 35, 31]. If these deviations also correspond to measurable shifts in brain activation, embedding-based models could offer novel insights into brain-language relationships and help inform computational mental health diagnostics.
Beyond clinical contexts, this framework could further support advanced applications in areas such as LLM-generated text detection [36], where subtle differences in linguistic patterns, coherence, and semantic clustering could indicate non-human authorship.
This study, therefore, investigates whether it is possible to computationally map natural language emotional expressions directly to brain regions without neuroimaging data, using semantic embeddings and clustering techniques to bridge text analysis with neuro-anatomical mapping. Specifically, we aim to:
@@
    Develop a novel computational framework that transforms textual emotional content into neuro-anatomically plausible brain region activations using state-of-the-art natural language processing embedding techniques.
@@
    Validate the clinical utility of this approach by demonstrating its ability to differentiate between healthy and depressed populations through distinct emotional brain mapping patterns, building on established neuroimaging differences in depression.
@@
    Demonstrate discrete emotion localization by mapping specific emotional states to neuro-anatomically appropriate brain regions with high spatial specificity, consistent with established emotion-brain mapping literature.
  • 4.
    Provide an objective, brain-based measure of human-likeness in language by comparing LLM-generated and human-authored texts through their inferred emotional brain activation patterns.
By providing a cost-effective, scalable alternative to traditional neuroimaging methods, such a computational approach can open new avenues for understanding the neural basis of human emotional communication in digital environments, while maintaining the ability to generate interpretable, neuro-anatomically grounded predictions from textual data alone.
@@ 2.1 Datasets
Three text-based datasets were employed in this study (Table 1). The DIAC-WOZ dataset [37] comprises annotated interview transcripts from individuals diagnosed with depression and healthy controls. The GoEmotions dataset [38] includes 58,000 Reddit [39] comments manually labeled into 27 emotion categories (or neutral). The Schema-Guided Dialogue dataset [40] represents nearly half a million sentences comprised of human and LLM chatbot interactions. All datasets consist of texts produced by native English speakers.
@@
DAICWOZ[37] Healthy and Depressed Categories 134 Clinical interview transcripts
GoEmotions [38] 27 Emotion Categories 58k English Reddit comments
The Schema-Guided Dialogue Dataset [40] Human and Chat bot conversations 463,282 English sentences
@@ 2.2 Text Preprocessing and Embedding Generation
During preprocessing (Figure 1, step 1), text was segmented into chunks of approximately 300 characters using periods as sentence boundaries. Next, text embeddings were generated using OpenAI’s [41] text-embedding-ada-002 model (Figure 1, step 2), which produces 1536-dimensional vector representations of input text. This model was chosen for its strong performance in capturing both semantic relationships and emotional nuances in natural language [29, 30, 31, 32]. Although training a custom embeddings model for our experimentation is technically feasible, it could introduce unwanted bias into our experiments, while the OpenAI text-embedding-ada-002 embeddings are commonly used in many public and commercial LLMs. The programming pseudo-code for steps 1 and 2 of Fig. 1 is shown in Algorithm 1.
Figure 1: Five-step computational pipeline to convert natural language text to embeddings, reduce dimensionality, cluster to emotional groups and map to brain regions.
@@ 2.3 Dimensionality Reduction and Spatial Mapping
The high-dimensional embeddings underwent a dimensionality reduction process (Figure 1, step 3A) using Principal Component Analysis (PCA) to reduce the dimensionality to three components, representing the minimum number of dimensions required for spatial brain mapping.
@@ 2.4 Emotional Intensity Estimation
Emotional intensity was quantified using a lexicon-based approach [42, 43, 44, 45, 46, 47] combined with syntactic feature analysis (Figure 1, step 3B). This emotion lexicon was constructed containing multiple terms across intensity levels:
  • 1.
    Extreme intensity (1.0): Terms indicating maximum emotional activation (e.g., ”devastated,” ”euphoric,” ”depressed,” ”extremely”)
  • 2.
    High intensity (0.8): Strong emotional indicators (e.g., ”amazing,” ”hate,” ”terrible”)
  • 3.
    Moderate intensity (0.6): Common emotional expressions (e.g., ”love,” ”sad,” ”happy”)
  • 4.
    Mild intensity (0.3): Subtle emotional content (e.g., ”nice,” ”bad,” ”okay”)
Additional intensity modifiers were incorporated to capture amplification effects:
  • 1.
    Intensification modifiers (”so,” ”very,” ”really,” ”truly,” ”completely,” ”totally”) added 0.3 to base intensity
  • 2.
    Absolutist terms (”never,” ”always,” ”everything,” ”nothing”) contributed 0.2 additional intensity
  • 3.
    Exclamation marks added 0.25 per occurrence (maximum 4)
  • 4.
    Question marks contributed 0.15 per occurrence (maximum 3)
  • 5.
    All-caps text (more than 3 characters) added 0.5 intensity
Base intensity was set at 0.1 for all texts to account for implicit emotional content, with final intensity scores capped at 2.0 to prevent extreme outliers from skewing subsequent analyses. The programming pseudo-code for steps 3A and 3B of Fig. 1 is shown in Algorithm 2.
@@ 2.5 Emotion Region Clustering
K-means clustering was applied to the 3D PCA-transformed embeddings to identify distinct emotional patterns within the data (Figure 1, step 4). The number of clusters was set to match 29 predefined anatomical brain regions (Table 2), establishing a direct correspondence between emotional content clusters and neuro-anatomical structures [5, 1, 48, 49, 50]. Of the 29 anatomically defined brain regions selected, 14 have been consistently implicated in emotion processing [3, 51, 2, 1].
Table 2: Predefined anatomical brain regions (29) used for k-means clustering.
@@
Limbic System Amygdala Left, Amygdala Right, Hippocampus Left, Hippocampus Right,
Anterior Cingulate Left, Anterior Cingulate Right
Cortical Regions Prefrontal Cortex Left, Prefrontal Cortex Right, Medial Prefrontal Cortex,
Orbitofrontal Left, Orbitofrontal Right, Insula Left, Insula Right,
Temporal Pole Left, Temporal Pole Right, Superior Temporal Left,
Superior Temporal Right, Posterior Cingulate
Subcortical Regions Caudate Left, Caudate Right, Putamen Left, Putamen Right,
Nucleus Accumbens Left, Nucleus Accumbens Right, Hypothalamus,
Periaqueductal Gray, Ventral Tegmental Area, Raphe Nuclei,
Locus Coeruleus
2.6 Cluster-to-Region Neuroanatomical Mapping
The assignment of emotional content clusters to specific brain regions (Figure 1, step 5) employed a two-stage distance minimization approach that preserved the one-to-one mapping between clusters and anatomical brain regions.
First, each cluster center’s coordinates in the 3D PCA space were compared against the predefined Montreal Neurological Institute (MNI) [52] coordinate positions of the 29 emotion processing brain regions using Euclidean distance. These regions were selected to span key nodes of the emotional circuitry and included limbic structures, prefrontal regions, subcortical structures, temporal regions, and brain stem nuclei, along with anterior and posterior cingulate cortex and medial prefrontal cortex.
The mapping process utilized a greedy assignment algorithm where cluster centers were sequentially matched to their nearest available brain regions. For each cluster center, distances to all anatomical regions were calculated and sorted in ascending order. The cluster was then assigned to the closest region that had not yet been claimed by another cluster, with this constraint preventing multiple clusters from mapping to the same anatomical location.
This sequential assignment continued until all clusters were mapped to unique brain regions, ensuring that the spatial distribution of emotional content in the 3D space corresponded meaningfully to the anatomical organization of emotion-processing brain networks. Once cluster-to-region assignments were established, individual text samples inherited their brain region labels based on their cluster membership, creating the final mapping from textual content to neuro-anatomical locations. Figure 2 provides a visual explanation of this cluster to brain region mapping process.
@@ 2.6 Cluster-to-Region Neuroanatomical Mapping
The programming pseudo-code for steps 4 and 5 of Fig. 1 is shown in Algorithm 3. Figure 3 provides a visual hierarchy detailing the emotion to brain region mapping approach as detailed in steps 4 and 5 of Figure 1.
@@
20:   for idxidx in argsort(distancesdistancesdo

### v3

@@ Abstract
Understanding how emotional expression in language relates to brain function remains a key challenge in neuroscience. Traditional neuroimaging provides valuable insight but is costly and limited to controlled laboratory settings. Here, we present a computational framework that explores potential links between emotional content in natural language to neuro-anatomical regions associated with affective processing. This, when validated through complementary neuroimaging, may enable scalable, imaging-free investigation of emotion-brain relationships. Our approach combines text embeddings, dimensionality reduction, and clustering to identify emotional states, which are then mapped to relevant brain regions. The framework was evaluated across three applications: (i) comparing healthy and depressed individuals, (ii) analyzing a large-scale emotion dataset, and (iii) contrasting human and large language model (LLM) outputs. Emotion intensity was quantified using a lexical scoring system sensitive to keywords, syntax, and modifiers, producing computationally plausible emotion-to-region clusters with visualization mapping. Across experiments, the framework distinguished healthy from depressed participants through distinct computational activation patterns and revealed systematic differences between human and LLM-generated texts in predicted computational engagement. A key finding emerged: depressed individuals exhibited reduced emotional diversity, showing 2.2 - 2.7 times more homogeneous emotional expression than healthy controls, suggesting that emotional rigidity may serve as a computational marker of depression. These computational patterns represent testable hypotheses, requiring further validation through neuroimaging. This work establishes a scalable, cost-effective tool for advancing both clinical and computational models of emotion, and provides a neuro-inspired benchmark for assessing how closely AI-generated language mirrors human emotional expression.
@@
The integration of embeddings and neuro-anatomical mapping successfully differentiated between healthy and depressed populations through distinct activation patterns.
@@
Regional assignment patterns were derived from established neuroimaging coordinates, creating computationally-predicted activation patterns that require independent validation.
In favor of reproducible research and to advance the field, all programming code used in this study is made publicly available.
@@ 1 Introduction
Understanding the neural correlates of emotion has traditionally relied on neuroimaging modalities such as electroencephalography (EEG) and functional magnetic resonance imaging (fMRI) lindquist2012 , vytal2010 , murphy2003 . These approaches have identified key regions such as the amygdala, insula, anterior cingulate cortex, and prefrontal cortex as central to emotional processing saarimaki2022 , phillips2003 , sliz2012 , ibrahim2022 . However, neuroimaging studies face substantial challenges including high cost, limited accessibility, controlled laboratory constraints, and reduced ecological validity when studying naturalistic emotion drevets2001 , hastings2004 . Consequently, there is growing interest in computational alternatives capable of mapping emotion-brain relationships using naturalistic data sources such as text.
Recent advances in language modeling have shown striking parallels between large language model (LLM) embeddings and human brain activity. Caucheteux et al. caucheteux2022 demonstrated that pre-trained language models align with neural responses without task-specific training, while Toneva et al. toneva2022 and Schrimpf et al. schrimpf2021 confirmed geometric and predictive correspondence between LLM-derived representations and cortical activation patterns during language comprehension. These findings suggest that the distributed semantic representations learned by LLMs may approximate aspects of neural language encoding in a computational sense, providing a promising foundation for computationally modeling potential brain–language relationships. Similar to recent advances in visual recognition that emphasize non-parametric, embedding-based reasoning, our approach relies on representational geometry rather than parametric classification. The Deep Nearest Centroids (DNC) framework Wang2022DeepCentroids , for instance, achieves interpretable decision-making by associating test samples with class sub-centroids in embedding space, enabling both explainability and cross-domain transferability. This methodological alignment underscores the potential of embedding-based representations for interpretable, model-derived approximations of language-brain relationships.
Parallel work has explored how emotionally-charged language engages distinct neural systems. Tomasino et al. tomasino2023 and Chen et al. chen2023 showed that linguistic valence correlates with differential activation of prefrontal and limbic regions, aligning with meta-analytic findings that positive and negative emotions recruit left and right hemispheric structures, respectively davidson2004 . Zhou et al. zhou2022 and Xiao et al. xiao2021 further demonstrated that embeddings derived from emotional text can be linked to fMRI and EEG signals, highlighting distributed yet consistent mappings across emotion categories. Together, these studies underscore that emotional semantics in language may provide a viable proxy for corresponding neural patterns.
Beyond theoretical mapping, embedding-based models may capture clinically relevant differences in emotional and linguistic processing. Individuals with depression and related conditions exhibit altered word choice, affective tone, and syntactic complexity campbell2004 , brosch2022 , liu2025 , ge2025 , lorenzoni2025 , zhong2025 . If such language deviations correspond to altered brain activation patterns, computational emotion-brain mapping could enable scalable biomarkers for mental health ramirez2022 , schwartz2014 , zhou2021 . Relatedly, these methods may help distinguish between human and machine-generated text through their inferred emotional brain activation patterns bulat2023 .
Despite prior progress, there is no existing imaging-free framework that directly links natural language emotion to specific neuro-anatomical regions. The present study introduces a computational approach that transforms emotional embeddings into brain-region activation patterns. Specifically, we aim to:
@@
    Develop a novel computational framework for mapping emotional language to brain region coordinates derived from neuroimaging literature.
@@
    Generate testable hypotheses by applying this framework to differentiate between healthy and depressed populations.
@@
    Produce emotion-region association patterns as predictions requiring orthogonal validation through independent neuroimaging studies.
This computational approach offers a scalable, cost-effective alternative to neuroimaging, enabling interpretable emotion–brain mapping from text alone.
@@ 2.1 Datasets
Three text-based datasets were employed in this study (Table 1). The DIAC-WOZ dataset dataset_diac comprises annotated interview transcripts from individuals diagnosed with depression and healthy controls. The GoEmotions dataset dataset_emotions includes 58,000 Reddit reddit comments manually labeled into 27 emotion categories (or neutral). The Schema-Guided Dialogue dataset dataset_google represents nearly half a million sentences comprised of human and LLM chatbot interactions. All datasets consist of texts produced by native English speakers.
@@
DAICWOZdataset_diac Healthy and Depressed Categories 134 Clinical interview transcripts
GoEmotions dataset_emotions 27 Emotion Categories 58k English Reddit comments
The Schema-Guided Dialogue Dataset dataset_google Human and Chat bot conversations 463,282 English sentences
@@ 2.2 Text Preprocessing and Embedding Generation
Texts were divided into 300 character segments using sentence boundaries. Each segment was converted into a 1,536-dimensional vector using OpenAI’s text-embedding-ada-002 model openai (Figure 1, Step 2). This model was selected for its well-validated emotional and semantic coverage liu2025 , ge2025 , lorenzoni2025 , zhong2025 , avoiding bias that might arise from custom embeddings. The programming pseudo-code for steps 1 and 2 of Fig. 1 is shown in Algorithm 1.
Figure 1: Five-step computational pipeline to convert natural language text to embeddings, reduce dimensionality, cluster to emotional groups, map to brain regions and calculate activations.
@@ 2.3 Dimensionality Reduction and Spatial Mapping
The high-dimensional embeddings underwent a dimensionality reduction process (Figure 1, step 3) using Principal Component Analysis (PCA) to reduce the dimensionality to three components, representing the minimum number of dimensions required for spatial brain mapping. As PCA was used primarily to obtain a spatially interpretable representation for brain-region visualization, reducing embeddings to three components captures only a small portion of the total variance. Clustering performed within this 3D space therefore reflects an intentional interpretability trade-off rather than an assumption that these components preserve most of the embedding structure. The choice of a 3D subspace was made to enable direct mapping onto MNI coordinates and to support the cortical surface visualizations, while acknowledging that clustering in a low-variance space may limit the capture of finer-grained embedding structure.
@@ 2.4 Emotional Intensity Estimation
Emotional intensity was computed through a lexicon-based scoring scheme combined with syntactic modifiers (Figure 1, Step 3). Words were assigned base intensities (mild 0.3, moderate 0.6, high 0.8, extreme 1.0), adjusted by amplifiers (very, really 0.3), absolutists (always, never 0.2), and punctuation cues (0.25, 0.15). Uppercase text added 0.5; all scores were capped at 2.0. This weighting follows continuous affect-intensity principles from established lexica (NRC mohammad2018word , ANEW bradley1999anew ) rather than empirically tuned parameters. Algorithm 2 outlines the combined dimensionality reduction, spatial mapping and intensity estimation steps denoted in Figure 1 as Step 3.
@@ 2.5 Emotion Region Clustering
K-means clustering was applied to the 3D PCA-transformed embeddings to identify distinct emotional patterns within the data (Figure 1, step 4). The number of clusters was set to match 29 predefined anatomical brain regions, establishing a direct correspondence between emotional content clusters and neuro-anatomical structures phillips2003 , lindquist2012 , kober2008 , etkin2011 , seeley2007 . Of the 29 anatomically defined brain regions selected, 14 (Table 2) have been consistently implicated in emotion processing murphy2003 , phan2002 , vytal2010 , lindquist2012 .
Table 2: Predefined anatomical brain regions used for K-means clustering.
@@
Frontal Lobe Medial Orbitofrontal (bilateral), Lateral Orbitofrontal (bilateral),
Pars Opercularis (bilateral), Rostral Middle Frontal (bilateral), Superior Frontal (bilateral)
Temporal Lobe Parahippocampal (bilateral), Fusiform (bilateral), Entorhinal (bilateral)
Cingulate Gyrus Rostral Anterior Cingulate (bilateral), Caudal Anterior Cingulate (bilateral), Posterior Cingulate (bilateral)
Insula Insula (bilateral)
Occipital Lobe Lingual (bilateral), Cuneus (bilateral)
2.6 Cluster-to-Region Neuro-anatomical Mapping
Cluster centroids were matched to Montreal Neurological Institute (MNI) MNI2025 coordinates of the 29 target regions using Euclidean-distance minimization, enforcing a one-to-one mapping (Figures 2 and 3). Each text segment inherited the brain-region label of its assigned cluster. This mapping leverages published MNI coordinates as anatomical constraints; therefore, resulting correspondences reflect computational predictions rather than independent neuroimaging validation. Algorithm 3 details the assignment procedure. Figure 2 provides a visual explanation of this cluster to brain region mapping process.
@@ 2.6 Cluster-to-Region Neuro-anatomical Mapping
20:   for idxidx in argsort(distancesdistances) do